5. We wish to conduct a hypothesis test of the form Ho : μ1-2-0 vs Ha : μί-μ2 > 0. Both populations are assumed normal with equal variance and samples are assumed independent. We draw 15 observations for the first sample and 17 observations for the second sample. State the distribution of the test statistic for this test. (a) t-distribution with 32 degrees of freedom (b) t-distribution with 30 degrees of freedom (e) Normal distribution with 30 degrees of freedom (d) t-distribution with 31 degrees of freedom
Answer #1

We do not know the population standard deviations so this will be a t test. Also,

Degrees of freedom = n1 + n2 - 2 = 15 + 17 - 2 = 30

Hence,

Option B is correct.